Problem: Given information about a circle, rectangle and a triangle, calculate the area of the shape from the information supplied. Write pseudocode to solve this problem and write a Python program that asks the user to input the required information for a shape and then calculates the area of the shape.

Geometry Calculator
1.  Calculate the Area of a Circle
2.  Calculate the Area of a Rectangle
3.  Calculate the Area of a Triangle
4.  Quit

Enter your choice (1 - 4):
  • If the user enters 1, the program should ask for the radius of the circle and then display its area.
  • If the user enters 2, the program should ask for the length and width of the rectangle and then display the rectangle's area.
  • If the user enters 3, the program should ask for the length of the triangle's base and its height, and then display its area.
  • If the user enters 4, the program should end.

Input Validation:

  • Do not accept values outside the range of 1 and 4 when selecting from the menu.
  • Do not accept negative or less than 0 values for any input amounts (ie. radius, length, width, etc).

Can anybody help me with this? I want to get a Psuedocode for the problem and the Program code for this.
Answer #1

def ch(a):
if a == 1:
r= float(input('Enter the radius:'))
area =3.14*r*r
print('The area of the circle is %0.2f' %area)
elif a == 2:
l= float(input('Enter the length:'))
b= float(input('Enter the breadth:'))
area =l*b
print('The area of the rectangle is %0.2f' %area)
elif a == 3:
h= float(input('Enter the height:'))
b= float(input('Enter the base:'))
area =0.5*h*b
print('The area of the triangle is %0.2f' %area)
else:
print (exit)

print('Geometry Calculator \n')
print('1 Calculate the Area of Circle \n')
print('2 Calculate the Area of Rectangle \n')
print('3 Calculate the Area of Triangle \n')
print('4 Quit\n')
a= int(input('Enter your Choice:'))
ch(a)